How many diagonals and triangles can be drawn far a polygon having 17 congruent sides?
Answer by Edwin McCravy(20055)

Suppose the polygon is ABCDEFGHIJKLMNOPQ

Then each choice of 2 letters represents either a side or a diagonal.

So there are C(17,2) sides and diagonals altogether.

There are 17 sides, AB, BC, CD, ..., PQ so the number of diagonals is
17 less than C(17,2).

Answer:  C(17,2) - 17 = 136 - 17 = 119 diagonals.

Edwin